OpenStudy (anonymous):

Austin wants to sketch the graph of a circle represented by the given equation. (x + 6)2 + (y + 2)2 = 16 Using complete sentences, describe the steps used to sketch the circle on a coordinate grid

OpenStudy (anonymous):

@jim_thompson5910 @Hero

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

can you identify the center?

OpenStudy (anonymous):

It would be (-6,-2)?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

the radius is____

OpenStudy (anonymous):

8 right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

in general, the radius of (x-h)^2+(y-k)^2=r^2 is r

jimthompson5910 (jim_thompson5910):

so for example, the radius of (x-2)^2+(y-9)^2=7^2 is r = 7

OpenStudy (anonymous):

Alrighty

jimthompson5910 (jim_thompson5910):

so the radius of (x + 6)^2 + (y + 2)^2 = 16 is what?

OpenStudy (anonymous):

I wanna say -8 but i know thats wrong, so idk, im not good at this im sorry

jimthompson5910 (jim_thompson5910):

its ok, just don't give up

jimthompson5910 (jim_thompson5910):

16 is the same as 4^2, agreed?

OpenStudy (anonymous):

oh yes, agree

jimthompson5910 (jim_thompson5910):

ok so that matches with the r^2

jimthompson5910 (jim_thompson5910):

so basically r^2 = 16 r^2 =4^2 r = ???

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so to sum up what you found so far center: (-6,-2) radius: 4

jimthompson5910 (jim_thompson5910):

those two combined together uniquely determine a circle

jimthompson5910 (jim_thompson5910):

so basically that's all you need to draw the circle

jimthompson5910 (jim_thompson5910):

if you were doing it on paper, you would start with a blank xy axis like this 

jimthompson5910 (jim_thompson5910):

you would then plot the center (-6,-2)